Masterson Company's budgeted production calls for 71,000 liters in April and 67,000 liters in May of a key raw material that costs $1.85 per liter. Each month's ending raw materials inventory should equal 25% of the following month's budgeted materials. The April 1 inventory for this material is 17,750 liters. What is the budgeted materials need in liters for April?
Answer #1

Solution:

Ending materials inventory for April = 67000*25% = 16,750 liters

Budgeted materials needs for April = 71000 + 16750 = 87,750 liters